I need to calculate the proportion of ultrasound that passes from a transducer into soft tissue if the matching layer thickness is 0.072mm
The transducer is a 4MHz with a z=3x10^7 and c= 4000m/s
And the matching layer has a z=1.8x10^6kg.m.s and a density of 1159kg.m
1
Expert's answer
2020-03-11T14:18:13-0400

The ratio of the intensity of the reflected wave over the incident (transmitted) wave is


"a=\frac{(Z_2-Z_1)^2}{(Z_1+Z_2)^2}=\frac{(3\cdot10^7-1.8\cdot10^6)^2}{(1.8\cdot10^6+3\cdot10^7)^2}=0.786."